# Long-VITA: Scaling Large Multi-modal Models to 1 Million Tokens with Leading Short-Context Accuracy

### v1

@@ Abstract
Establishing the long-context capability of large vision-language models is crucial for video understanding, high-resolution image understanding, multi-modal agents and reasoning. We introduce Long-VITA, a simple yet effective large multi-modal model for long-context visual-language understanding tasks. It is adept at concurrently processing and analyzing modalities of image, video, and text over 4K frames or 1M tokens while delivering advanced performances on short-context multi-modal tasks. We propose an effective multi-modal training schema that starts with large language models and proceeds through vision-language alignment, general knowledge learning, and two sequential stages of long-sequence fine-tuning. We further implement context-parallelism distributed inference and logits-masked language modeling head to scale Long-VITA to infinitely long inputs of images and texts during model inference. Regarding training data, Long-VITA is built on a mix of $17$M samples from public datasets only and demonstrates the state-of-the-art performance on various multi-modal benchmarks, compared against recent cutting-edge models with internal data. Long-VITA is fully reproducible and supports both NPU and GPU platforms for training and testing. We hope Long-VITA can serve as a competitive baseline and offer valuable insights for the open-source community in advancing long-context multi-modal understanding.
@@ 3.2 Data Construction
Name # Param. Device Type Devices Number Inference Training 8 16 32 64 LongVILA [11] 7B 80G GPU $276$K $666$K Long-VITA 14B 64G NPU $1024$K $1024$K 96G GPU $400$K $800$k $1600$K $1024$K

### v2

@@ Abstract
We introduce Long-VITA, a simple yet effective large multi-modal model for long-context visual-language understanding tasks. It is adept at concurrently processing and analyzing modalities of image, video, and text over 4K frames or 1M tokens while delivering advanced performances on short-context multi-modal tasks. We propose an effective multi-modal training schema that starts with large language models and proceeds through vision-language alignment, general knowledge learning, and two sequential stages of long-sequence fine-tuning. We further implement context-parallelism distributed inference and logits-masked language modeling head to scale Long-VITA to infinitely long inputs of images and texts during model inference. Regarding training data, Long-VITA is built on a mix of $17$M samples from public datasets only and demonstrates the state-of-the-art performance on various multi-modal benchmarks, compared against recent cutting-edge models with internal data. Long-VITA is fully reproducible and supports both NPU and GPU platforms for training and testing. By leveraging our inference designs, Long-VITA models achieve a remarkable 2$\times$ prefill speedup and 4$\times$ context length extension in single node with $8$ GPUs. We hope Long-VITA can serve as a competitive baseline and offer valuable insights for the open-source community in advancing long-context multi-modal understanding.
@@ 3.2 Data Construction
Name # Param. Device Type Devices Number Inference Training 8 16 32 64 LongVILA [11] 7B 80G GPU 276K  [^6] K Long-VITA 14B 64G NPU 1,024K 1,024K 96G GPU 400K 800K 1,600K 1,024K
Table 4: Comparison of different language modeling head.
# GPUs Method # Frames # Tokens Prefill Time (s) 8 Original LM Head 390 100K 19.2 400 103K OOM Logits-Masked LM Head 390 100K 10.1 ( $\downarrow$ 47.3% ) 1,620 417K ( $\uparrow$ 417% ) 102 1,630 420K OOM 32 Chunked LM Head 4,096 1,056K 177 6,400 1,651K 377 Logits-Masked LM Head 4,096 1,056K 157 ( $\downarrow$ 11.3% ) 6,400 1,651K 335 ( $\downarrow$ 11.1% )
@@ 3.4 Hyper-parameters and Infrastructures
Tab. 4 further shows the effectiveness of logits-masked language modeling head (logits-masked LM head). All methods are implement with Flash Attention and context-parallelism distributed inference on GPU with $96$G memory and about $150$ TFLOPS for bfloat16. Compared to the original LM head, logits-masked LM head extends the max sequence length by $417\%$, and reduces time cost by $47.3\%$. We also implement a chunked language modeling head (chunked LM head), which process tokens with a chunk length of $32,768$. Compared to the chunked LM head, logits-masked LM head achieves $11.3$ and $11.1$ speedup under the $1$M and $1.6$M input lengths, respectively.